In a circular coil of 15.0 cm the current is 3.50 A. The coil is in a magnetic field of magnitude of 15.0 T in such a way that its normal vector makes an angle of 30 degrees with the field. Determine the coil torque and the magnetic dipole moment.
Answer #1

magnetic moment

\mu = NiA = 1 (3.50 A) (\pi*0.152) = 0.247 Am2

Torque

\tau = \mu B sin\theta =  0.247 Am2 * 15 T * sin 30 = 1.855 Nm